### v1

# Explaining higher-order correlations between elliptic and triangular flow

Abstract
The ALICE Collaboration has analyzed a number of cumulants mixing elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), involving up to $8$ particles, in Pb+Pb collisions at the LHC. We unravel an unexpected simplicity in these complex mathematical quantities for collisions at fixed impact parameter. We show that as one increases the order in $v_{2}$, for a given order in $v_{3}$, the changes in the cumulants are solely determined by the mean elliptic flow in the reaction plane, which originates from the almond-shaped geometry of the overlap area between the colliding nuclei. We derive simple analytic relations between cumulants of different orders on this basis. Some of these relations are in reasonable agreement with existing data. We postulate that agreement will be much improved if the analysis is repeated with a finer centrality binning and a larger pseudorapidity acceptance.
I Introduction
Cumulants of multiparticle azimuthal correlations are routinely analyzed in nucleus-nucleus collisions STAR:2002hbo ; NA49:2003njx ; ALICE:2010suc ; ALICE:2011ab ; CMS:2012zex ; CMS:2013wjq ; ATLAS:2014qxy ; STAR:2015mki ; ALICE:2018lao . They were originally introduced Borghini:2000sa ; Borghini:2001vi in order to separate elliptic flow in the direction of impact parameter, induced by pressure gradients within the almond-shaped quark-gluon plasma Ollitrault:1992bk , from other correlations of various origins (Bose-Einstein Dinh:1999mn , global momentum conservation Borghini:2000cm ; Bzdak:2010fd , resonance decays Feng:2021pgf , jets Kovchegov:2002nf ; Connors:2017ptx ), referred to as “nonflow” correlations, whose contribution decreases rapidly as the order of the cumulant (the number of correlated particles) increases. The perspective changed when it was realized that local fluctuations in the initial energy density contribute significantly to pressure gradients, generating large event-by-event fluctuations of elliptic flow Miller:2003kd ; Andrade:2006yh ; PHOBOS:2006dbo , and a new phenomenon solely due to fluctuations, triangular flow Alver:2010gr .
This led to a paradigm shift. One now considers that flow is the dominant contribution to all cumulants, even the lowest-order ones (pair correlations), provided that there is a rapidity gap between the particles that are correlated PHENIX:2003qra . Thus, cumulants are solely determined by the probability distribution of flow fluctuations ATLAS:2013xzf ; ALICE:2018rtz , which is used in order to constrain models of the initial energy density Retinskaya:2013gca ; Moreland:2014oya . This turns out to be the most robust aspect of global theory-to-data comparison JETSCAPE:2020mzn .
In this paper, we unravel the information contained in cumulants of azimuthal correlations involving both elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), which are called “mixed harmonic cumulants” (MHC) ALICE:2021adw . $v_{2}$ and $v_{3}$ are the largest Fourier harmonics of anisotropic flow ALICE:2011ab . Relativistic hydrodynamics predicts Niemi:2012aj that they are approximately proportional to the initial anisotropies $\varepsilon_{2}$ and $\varepsilon_{3}$, which are Fourier coefficients of the initial density profile. Therefore, cumulants involving $v_{2}$ and $v_{3}$ constrain models of the initial state more directly than higher harmonics, which are largely driven by nonlinear response terms Borghini:2005kd ; Gardim:2011xv . The lowest-order MHC is a 4-particle cumulant introduced by Bilandzic et al. Bilandzic:2013kga , which has been measured in Pb+Pb collisions at the LHC by ALICE ALICE:2016kpq and ATLAS ATLAS:2019peb . It represents the linear correlation between $v_{2}^{2}$ and $v_{3}^{2}$, which is negative except in ultra-central collisions. We have recently shown Alqahtani:2025wan that it is driven by the correlation between the elliptic flow in the reaction plane, $v_{2,x}$, and $v_{3}^{2}$. Our goal is to extend this study to higher-order cumulants, of 6 and 8 particles, which have subsequently been measured ALICE:2021adw .
Throughout this paper, we assume that nonflow correlations are negligible, so that particles in each event are emitted independently according to an underlying probability distribution Luzum:2011mm ; Ollitrault:2023wjk . Let $P(\varphi)$ denote the azimuthal dependence of this probability distribution. The complex anisotropic flow $V_{n}$ of the event is defined as its Fourier coefficient of order $n$ Luzum:2013yya :
$V_{n}\equiv\int_{0}^{2\pi}e^{in\varphi}P(\varphi)d\varphi.$ (1)
The usual anisotropic flow $v_{n}$ is the modulus, $v_{n}\equiv|V_{n}|$.
In Sec. II, we define the cumulants of $v_{2}$ and $v_{3}$ measured by ALICE using the formalism of generating functions and we briefly discuss the data. In Sec. III, we define similar cumulants but in a different coordinate frame, the “intrinsic” frame where the direction of impact parameter is fixed Roubertie:2025qps ; Alqahtani:2025wan . In this frame, the magnitude of cumulants decreases as a function of the order, in a way which follows specific scaling rules, provided that all events have the same impact parameter. In Sec. IV, we express the experimental cumulants as a function of the cumulants in the intrinsic frame. We explain the orders of magnitude of cumulants measured by ALICE. In Sec. V, we derive relations between cumulants of different orders, and compare them with experimental data.
II Cumulants of flow fluctuations in the laboratory frame
The quantities that can be measured experimentally are moments of the joint distribution of $V_{2}$ and $V_{3}$ Bhalerao:2011yg ; Bhalerao:2014xra . The cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$ analyzed by ALICE ALICE:2021adw , where $m$ and $q$ are positive integers, are combinations of these moments Bilandzic:2010jr ; Bilandzic:2013kga . [^1] The ALICE collaboration has subsequently analyzed a different set of “cumulants” which are different combination of moments ALICE:2023lwx . They are not derived from the same generating function as the cumulants studied in this paper, and our analysis does not apply to them. Their expressions are derived by expanding the generating function of cumulants Taghavi:2020gcy , which depends on complex variables Mehrabpour:2020wlu $\lambda$ and $\mu$ and their complex conjugates $\lambda^{*}$ and $\mu^{*}$:
$G_{\rm lab.}(\lambda,\mu)$ $\equiv\ln\left\langle\exp\left(\lambda^{*}V_{2}+\lambda V_{2}^{*}+\mu^{*}V_{3}+\mu V_{3}^{*}\right)\right\rangle$
$=\sum_{m,q\geq 0}\frac{(\lambda\lambda^{*})^{m}(\mu\mu^{*})^{q}}{(m!)^{2}(q!)^{2}}MHC(v_{2}^{2m},v_{3}^{2q}).$ (2)
In the first line, angular brackets in the right-hand side denote an average over events in a centrality class. In the second line, we have only kept the relevant terms in the power-series expansion. [^2] Azimuthal symmetry requires that the generating function is symmetric under $(\lambda,\mu)\to(\lambda e^{2i\varphi},\mu e^{3i\varphi})$. This cancels many terms in the power series expansion, but terms involving $\lambda^{3}(\mu^{*})^{2}$ are also non-vanishing (see Sec. III.1). We omit them for simplicity since we do not study them in this paper. They correspond to angular correlations between $V_{2}$ and $V_{3}$ Bhalerao:2011yg , also referred to as event-plane correlations ATLAS:2014ndd . $MHC(v_{2}^{2m},v_{3}^{2q})$ is a cumulant of order $2(m+q)$. If $q=0$ or $m=0$, it corresponds to the cumulant of a single harmonic $n=2$ or $n=3$, usually denoted by $c_{n}\{2k\}$ Borghini:2001vi ; ATLAS:2019peb (see Appendix A):
$MHC(v_{2}^{2m},v_{3}^{0})$ $=c_{2}\{2m\}$
$MHC(v_{2}^{0},v_{3}^{2q})$ $=c_{3}\{2q\}.$ (3)
The mixed cumulants are those for which both $m$ and $q$ are positive. The lowest-order mixed cumulant corresponds to $m=q=1$. Expanding the left-hand side of Eq. (2) to order $\lambda\lambda^{*}\mu\mu^{*}$, one obtains its expression in terms of moments:
$MHC(v_{2}^{2},v_{3}^{2})=\langle v_{2}^{2}v_{3}^{2}\rangle-\langle v_{2}^{2}\rangle\langle v_{3}^{2}\rangle.$ (4)
It was measured by ALICE in 2016 in Ref. ALICE:2016kpq , where it was named $SC(3,2)$. Higher-order cumulants with $(m,q)=(2,1),(3,1),(1,2),(2,2),(1,3)$ were subsequently measured in Ref. ALICE:2021adw , where their expressions in terms of moments are provided. Deriving these expressions is straightforward using Eq. (2). We do not repeat them because they are lengthy and will not be needed in this work. [^3] Note that the six MHC cumulants analyzed by ALICE only involve the magnitudes, not the phases of $V_{2}$ and $V_{3}$. The phases only enter at the next order $(m,q)=(3,2)$, a cumulant of order 10 whose expression involves moments of the type $\langle V_{2}^{3}(V_{3}^{*})^{2}\rangle$, and which has not yet been analyzed.
Figure 1: Comparison between ALICE ALICE:2021adw and ATLAS ATLAS:2019peb data on $nMHC(v_{2}^{2},v_{3}^{2})$ in Pb+Pb collisions at 5.02 TeV per nucleon pair, as a function of the collision centrality. ATLAS denotes this quantity by $nsc_{2,3}\{4\}$.
ALICE normalizes the cumulants as follows:
$nMHC(v_{2}^{2m},v_{3}^{2q})\equiv\frac{MHC(v_{2}^{2m},v_{3}^{2q})}{\langle v_{2}^{2m}\rangle\langle v_{3}^{2q}\rangle}.$ (5)
This normalization suppresses the sensitivity to kinematic cuts, and also provides an intuitive, dimensionless measure of the strength of the correlation. A value of order unity would imply a strong correlation between $v_{2}$ and $v_{3}$. But the measured values of $nMHC(v_{2}^{2m},v_{3}^{2q})$ are, in absolute magnitude, smaller than unity for all centralities. [^4] By contrast, $nMHC(v_{2}^{2},v_{4}^{2})$ reaches $1$ in peripheral collisions ATLAS:2019peb ; ALICE:2021adw , which is a natural consequence of the non-linear coupling between $v_{2}$ and $v_{4}$ Teaney:2012ke . They increase somewhat with $m$, and decrease strongly with $q$, which will be explained in Sec. IV.
The normalization (5) also facilitates the comparison with other experiments and theory calculations. A comparison between ALICE and ATLAS ATLAS:2019peb data for the lowest-order cumulant $nMHC(v_{2}^{2},v_{3}^{2})$ is displayed in Fig. 1. There are sizable differences between the two experiments, whose origin is unknown. One possible explanation is the wider centrality bins used by ALICE. However, one would typically expect wider bins to increase the value of $nMHC(v_{2}^{2},v_{3}^{2})$ Gardim:2016nrr , and ALICE is below ATLAS for most centralities. Interestingly, ATLAS observes a variation of $nMHC(v_{2}^{2},v_{3}^{2})$ going in the same direction as ALICE (down for centralities $40\%$, up for centralities $40\%$) when only particles with $p_{T}>2$ GeV/c/c are included. Since these high-$p_{T}$ particles are likely to have sizable nonflow correlations from jet production, it is tempting to postulate that the difference between ATLAS and ALICE may be due to larger nonflow effects in ALICE, due to the smaller acceptance in pseudorapidity. We consider it likely that ATLAS results are more precise than ALICE results, and we will take this as an excuse for not understanding precisely the ALICE results on higher-order cumulants in Sec. V.
Figure 2: Mixed harmonic cumulants as a function of centrality for Pb+Pb collisions at 5.02 TeV per nucleon pair. The normalized cumulants (5) are taken from Ref. ALICE:2021adw , and the moments in the denominator are evaluated using earlier ALICE data ALICE:2016ccg and formulas in Appendix A. For the negative cumulants, we change the sign before taking the logarithm. $MHC(v_{2}^{4},v_{3}^{4})$ and $MHC(v_{2}^{2},v_{3}^{6})$ are positive for the most central bins, and the three corresponding data points are circled.
For our analysis, we will need the un-normalized cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$. In order to compute them, we evaluate the moments appearing in the denominator of Eq. (5) using standard formulas which are recalled in Appendix A. Results are displayed in Fig. 2 as a function of the collision centrality. The lowest-order cumulant defined by Eq. (4) is negative, as shown in Fig. 1, and the sign alternates as $m$ or $q$ increases, as illustrated by the legend of the figure.
III Cumulants of flow fluctuations in the intrinsic frame
Figure 3: Schematic representation of two collision events with the same impact parameter seen in the laboratory frame and in the intrinsic frame. The dots in the overlap area correspond to the positions of participants nucleons at the time of impact, which takes a snapshot of the nuclear wavefunction.
We now carry out a thought experiment where all events are aligned in such a way that their impact parameter is along the $x$ axis. [^5] The $(x,z)$ plane is the reaction plane, and the $x$ axis is often called the reaction plane by a slight abuse of language. We call this the “intrinsic frame” Alqahtani:2024ejg ; Roubertie:2025qps , as illustrated in Fig. 3. More detailed information is available in the intrinsic frame than in the laboratory frame, in which the direction of impact parameter is unknown. Only a small subset of this information can be reconstructed from experimental data, as will be illustrated below. But this limited information sheds light on the physics underlying the measured cumulants.
III.1 Definitions and general properties
The cumulants of $V_{2}$ and $V_{3}$ in the intrinsic frame are
$c_{mpqr}\equiv\langle(V_{2})^{m}(V_{2}^{*})^{p}(V_{3})^{q}(V_{3}^{*})^{r}\rangle_{c},$ (6)
where the subscript $c$ means that one isolates the connected part of the correlation (the cumulant) by subtracting the contributions of lower-order correlations order by order.
For central collisions with $b=0$, azimuthal symmetry implies that $c_{mpqr}$ may differ from 0 only if 2(mp)+3(qr)=02(m-p)+3(q-r)=0. A sufficient condition is to have $m=p$ and $q=r$. The corresponding cumulants $c_{mmqq}$ coincide with $MHC(v_{2}^{2m},v_{3}^{2q})$, as will be shown below.
For non-central collisions, azimuthal symmetry is lost since the $x$ axis is the reaction plane. Two weaker symmetries remain: Symmetry with respect to the reaction plane, $\varphi\to-\varphi$ (Fig. 3), implies that $c_{mpqr}$ is real and $c_{pmrq}=c_{mpqr}$. Symmetry between target and projectile, $\varphi\to\varphi+\pi$, further implies that $c_{mpqr}$ vanishes for odd $q+r$.
The essential difference between the intrinsic frame and the laboratory frame is that there is now a non-trivial cumulant of order $1$ allowed by symmetry, namely, the mean elliptic flow in the reaction plane. We denote it by $\bar{V}_{2}$:
$\bar{V}_{2}\equiv c_{1000}=c_{0100}.$ (7)
There are four cumulants of order $2$ allowed by symmetry: $c_{1100}$, $c_{0011}$, $c_{0200}=c_{2000}$ and $c_{0002}=c_{0020}$. The first two are the largest and correspond to the width of $v_{n}$ fluctuations, $\sigma_{v_{n}}$ Voloshin:2007pc : [^6] The other two cumulants $c_{0200}$ and $c_{0002}$ measure the asymmetry of $v_{n}$ fluctuations, and are much smaller, as will be shown in Sec. III.2.
$\sigma_{v_{2}}^{2}$ $\equiv c_{1100}$
$\sigma_{v_{3}}^{2}$ $\equiv c_{0011}.$ (8)
The central limit theorem implies that only cumulants of order $1$ and $2$ are kept, which means that the probability distribution of $V_{2}$ and $V_{3}$ is Gaussian in the intrinsic frame Voloshin:2007pc . Since none of these cumulants mix $V_{2}$ and $V_{3}$, they are uncorrelated in the Gaussian limit Alqahtani:2025wan . Therefore, the correlations studied in this paper are all driven by non-Gaussian properties of the distribution, so that we shall not neglect cumulants of order $3$ and higher. [^7] Non-Gaussian fluctuations are also responsible for the non-zero $v_{3}\{4\}$ seen in Pb+Pb ALICE:2011ab and Xe+Xe CMS:2019cyz ; CMS:2025xtt collisions , the hierarchy of higher-order cumulants of $v_{2}$ ($v_{2}\{4\}$, $v_{2}\{6\}$, $v_{2}\{8\}$) in p+Pb collisions CMS:2015yux ; CMS:2019wiy , as well as the splitting between $v_{2}\{4\}$ and $v_{2}\{6\}$ in mid-central Pb+Pb collisions Giacalone:2016eyu ; CMS:2017glf ; ALICE:2018rtz ; CMS:2023bvg . In Sec. III.2, we discuss the orders of magnitude of these higher-order cumulants and, more specifically, how they depend on the size and shape of the quark-gluon plasma created in the early stages of the collision.
Cumulants defined by Eq. (6) are expressed as a function of moments using a generating function which is formally identical to Eq. (2), the only difference being that one averages over events with the same reaction plane:
$G_{\rm intr.}(\lambda,\mu)$ $\equiv\ln\left\langle\exp\left(\lambda^{*}V_{2}+\lambda V_{2}^{*}+\mu^{*}V_{3}+\mu V_{3}^{*}\right)\right\rangle$
$=\sum_{m,p,q,r\geq 0}\frac{(\lambda^{*})^{m}\lambda^{p}(\mu^{*})^{q}\mu^{r}}{m!\,p!\,q!\,r!}c_{mpqr}.$ (9)
For central collisions with $b=0$, azimuthal symmetry is restored, and the generating functions (2) and (9) are identical. This implies $c_{mmqq}=MHC(v_{2}^{2m},v_{3}^{2q})$. [^8] Note, however, that these cumulants represent only a subset of the non-vanishing cumulants, since all cumulants with 2(mp)+3(qr)=02(m-p)+3(q-r)=0 are allowed by symmetry. For $b\not=0$, expressions of $MHC(v_{2}^{2m},v_{3}^{2q})$ as a function of intrinsic cumulants $c_{mpqr}$ involves additional terms, which will be derived in Sec. IV.
III.2 Power-counting scheme
We now derive general scaling laws that apply to the cumulants in the intrinsic frame. For simplicity, we neglect centrality fluctuations, and we assume that all events in a centrality class have the same impact parameter. The ALICE experiment uses wide centrality bins where this approximation may not be good enough, but this can easily be improved in future analyses. Centrality fluctuations always matter in the 0–5% centrality window Alqahtani:2024ejg , but we have found that their effect on correlations between $v_{2}$ and $v_{3}$ is modest beyond 5% centrality Alqahtani:2025wan , provided that the analysis uses narrow centrality bins ATLAS:2019peb , and we hope that the same conclusion applies to the higher-order cumulants studied in this paper.
If the impact parameter is constant, different events in the intrinsic frame differ only by local fluctuations in the initial density profile, which are of quantum origin, as illustrated in Fig. 3. The resulting fluctuations of the complex anisotropies $\varepsilon_{2}$ and $\varepsilon_{3}$ Qiu:2011iv , which generate $V_{2}$ and $V_{3}$ through linear hydrodynamic response, originate from a large number $N$ of independent local fluctuations. [^9] NN is typically of the same order as the number of participant nucleons Miller:2007ri . For independent fluctuations, a cumulant of order $k$ varies with $N$ like $N^{1-k}$, where the order is the sum of indices $k=m+p+q+r$. The variances in Eq. (8) are cumulants of order $2$, and are therefore of order $1/N$. We denote by $\cal V$} the typical magnitude of event-by-event flow fluctuations, as measured by $\sigma_{v_{2}}$ and $\sigma_{v_{3}}$, which are comparable Roubertie:2025qps , that is, ${\cal V}\propto 1/\sqrt{N}$. Then, a cumulant of order $k$ is of order $N^{1-k}\propto{\cal V}^{2k-2}$.
We now discuss how cumulants depend on the shape of the system, measured by its mean deformation. We assume that azimuthal symmetry is mildly broken, $\bar{V}_{2}\ll 1$. More specifically, we assume that $\bar{V}_{2}$ is also of order $\cal V$}, [^10] In practice, $\bar{V}_{2}<\sigma_{v_{n}}$ for central collisions, and $\bar{V}_{2}>\sigma_{v_{n}}$ for mid-central collisions Roubertie:2025qps . which implies that we treat formally $v_{2}$ and $v_{3}$ as quantities of the same order $\cal V$}. The order of magnitude of $c_{mpqr}$ depends on the shape through the order of the Fourier coefficient it corresponds to, 2n=|2(mp)+3(qr)|2n=|2(m-p)+3(q-r)|. For mildly-broken azimuthal symmetry, a Fourier coefficient of order $2n$ is suppressed by a factor $(\bar{V}_{2})^{n}\propto{\cal V}^{n}$. For instance, the asymmetries of $v_{2}$ and $v_{3}$ fluctuations, $c_{0200}$ and $c_{0002}$, are suppressed by factors ${\cal V}^{2}$ and ${\cal V}^{3}$ with respect to $c_{1100}$ and $c_{0011}$.
Putting together the dependences on the size and shape, we generally expect
$c_{mpqr}\sim{\cal O}\left({\cal V}^{2(m+p+q+r-1)+\left|m-p+\frac{3}{2}(q-r)\right|}\right).$ (10)
This gives for instance $c_{1111}\sim{\cal O}\left({\cal V}^{6}\right)$. Therefore, in central collisions, where $MHC(v_{2}^{2},v_{3}^{2})=c_{1111}$, each of the moments in the right-hand side of Eq. (4) is of order ${\cal V}^{4}$, while the difference is of order ${\cal V}^{6}$, i.e., much smaller. This systematic expansion scheme will allow us to single out the dominant contributions to each of the MHCs.
IV Relations between experimental cumulants and intrinsic cumulants
We now relate the two sets of cumulants defined in Secs. II and III. In the laboratory frame, the orientation of impact parameter is uniformly distributed. Therefore, moments of $V_{2}$, $V_{3}$, $V_{2}^{*}$, $V_{3}^{*}$ in the laboratory frame are obtained from the corresponding moments in the intrinsic frame by averaging over all possible orientations of impact parameter. This implies the following relation between the generating functions (2) and (9Abbasi:2017ajp :
$G_{\rm lab.}(\lambda,\mu)=\ln\left(\int_{0}^{2\pi}\frac{d\varphi}{2\pi}\exp\left(G_{\rm intr.}(\lambda e^{2i\varphi},\mu e^{3i\varphi})\right)\right).$ (11)
Expanding to order $\lambda\lambda^{*}\mu\mu^{*}$, one obtains the following exact expression of the lowest-order mixed cumulant (4) as a function of the intrinsic cumulants Alqahtani:2025wan :
$MHC(v_{2}^{2},v_{3}^{2})$ $=c_{1000}c_{0111}+c_{0100}c_{1011}+c_{1111}$
$=2\bar{V}_{2}c_{0111}+c_{1111}$ (12)
where, in the last equality, we have introduced $\bar{V}_{2}$ defined by Eq. (7) and used the symmetry $c_{mpqr}=c_{pmrq}$. For central collisions, $\bar{V}_{2}=0$ and $MHC(v_{2}^{2},v_{3}^{2})$ coincides with $c_{1111}$, which is of order ${\cal V}^{6}$ as explained at the end of Sec. III.2. For non-central collisions, the first term in the right-hand side of Eq. (12) differs from zero, but is also of order ${\cal V}^{6}$ according to Eq. (10).
We now list the expressions of the higher-order cumulants measured by ALICE ALICE:2021adw , which we truncate by keeping only the leading terms in the power-series expansion:
$MHC(v_{2}^{4},v_{3}^{2})$ $=-4\,\bar{V}_{2}^{3}\,c_{0111}$
$MHC(v_{2}^{6},v_{3}^{2})$ $=24\,\bar{V}_{2}^{5}\,c_{0111}$
$MHC(v_{2}^{2},v_{3}^{4})$ $=2(2\,c_{0111}^{2}+\bar{V}_{2}\,c_{0122})+c_{1122}$
$MHC(v_{2}^{4},v_{3}^{4})$ $=-4\bar{V}_{2}^{2}(6c_{0111}^{2}+\bar{V}_{2}c_{0122})$
$MHC(v_{2}^{2},v_{3}^{6})$ $=2(9c_{0111}c_{0122}+\bar{V}_{2}c_{0133})+c_{1133}.$ (13)
They are respectively of order ${\cal V}^{8}$, ${\cal V}^{10}$, ${\cal V}^{10}$, ${\cal V}^{12}$, ${\cal V}^{14}$. More generally, $MHC(v_{2}^{2m},v_{3}^{2q})$ is of order ${\cal V}^{2m+4q}$. These orders of magnitude are reflected in the hierarchy observed in Fig. 2. In particular, they explain why $MHC(v_{2}^{6},v_{3}^{2})$ and $MHC(v_{2}^{2},v_{3}^{4})$ are of comparable magnitude, despite being cumulants of different orders (8 and 6 respectively). Since $v_{2}$ and $v_{3}$ are both of order $\cal V$}, the normalized symmetric cumulant (5) is of order ${\cal V}^{2q}$. This explains why the magnitude of $nMHC(v_{2}^{2m},v_{3}^{2q})$ decreases strongly as $q$ increases, as pointed out in Sec. II.
We finally provide leading-order expressions for the cumulants of order 10, which have not yet been analyzed, but could easily be measured:
$MHC(v_{2}^{8},v_{3}^{2})=$ $-264\,\bar{V}_{2}^{7}\,c_{0111}$
$MHC(v_{2}^{6},v_{3}^{4})=$  $\,24\,\bar{V}_{2}^{4}(10\,c_{0111}^{2}+\bar{V}_{2}\,c_{0122})$
$MHC(v_{2}^{4},v_{3}^{6})=$ $-144\,\bar{V}_{2}\,c_{0111}^{3}$
$-4\bar{V}_{2}^{2}(27c_{0111}c_{0122}+\bar{V}_{2}c_{0133}).$ (14)
A first comment on the expressions (12), (13) and (14) is that they only involve the mean elliptic flow in the reaction plane, $\bar{V}_{2}$, and mixed cumulants of order $\geq 3$, which quantify non-Gaussian fluctuations. A second remark is that for a given value of $q$, $MHC(v_{2}^{2},v_{3}^{2q})$ contains more terms than higher-order cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$ with $m\geq 2$, whose expressions are somewhat simpler. More specifically, the expressions for $m=1$ contain the cumulant $c_{11qq}$ (note that it is the only term which does not vanish for central collisions), which no longer appears for $m\geq 2$. For a given value of $q$, each of the remaining terms is multiplied by a factor proportional to $\bar{V}_{2}^{2}$ as $m$ increases. This can be related to the well-known fact that cumulants of $v_{2}$ of order 4 and higher only depend on $\bar{V}_{2}$ to leading order in $\cal V$}, as recalled in Appendix A. In addition, there are overall numerical factors which increase rapidly as a function of $m$ (22, 44, $24$ for $m=1,2,3$). These numerical factors explain why the normalized cumulants $nMHC(v_{2}^{2m},v_{3}^{2q})$ increase as a function of $m$ for fixed $q$, in absolute magnitude ALICE:2021adw .
V Relations between cumulants of different orders
Figure 4: Ratios in Eqs. (15) and (17). Symbols are ALICE data, where the mixed cumulants are taken from Ref. ALICE:2021adw and $v_{2}\{4\}$ from Ref. ALICE:2016ccg , as a function of the collision centrality in Pb+Pb collisions at 5.02 TeV per nucleon pair. Horizontal lines are our theory predictions.
Eqs. (12), (13) and (14) show that to leading order, the experimental cumulants only depend on a small number of intrinsic cumulants. First of all, they involve the mean elliptic flow in the reaction plane, which can be obtained from data using the approximation $\bar{V}_{2}\approx v_{2}\{4\}$ (App. A). Since $v_{2}\{4\}$ is only measured above 5% centrality ALICE:2016ccg , we exclude the most central bin 0-5% from now on.
The remaining intrinsic cumulants can be eliminated by combining the information from several measured cumulants. Taking the ratio between the first two lines of Eq. (13), for instance, one obtains:
$\frac{MHC(v_{2}^{6},v_{3}^{2})}{v_{2}\{4\}^{2}MHC(v_{2}^{4},v_{3}^{2})}=-6.$ (15)
This prediction is tested against ALICE data in Fig. 4. The experimental ratio in the left-hand side is in rough agreement with the predicted value in the right-hand side, but somewhat smaller in absolute magnitude.
Similarly, using Eqs. (13) and (14), we predict:
$\frac{MHC(v_{2}^{8},v_{3}^{2})}{v_{2}\{4\}^{2}MHC(v_{2}^{6},v_{3}^{2})}=-11.$ (16)
This could easily be checked experimentally, as increasing the order in $v_{2}$ does not significantly increase errors.
Eqs. (15) and (16) are rigorous mathematical results to leading order in $\cal V$}. They generalize the well-known identities $v_{2}\{4\}=v_{2}\{6\}=v_{2}\{8\}$ (Appendix A) to mixed cumulants. We therefore expect that their accuracy is comparable, at the percent level in Pb+Pb collisions at LHC energies CMS:2017glf ; ALICE:2018rtz ; CMS:2023bvg . The discrepancy between our prediction and data is an order of magnitude larger than expected in Fig. 4. We predict that agreement will be much improved if the analysis is repeated with finer centrality bins.
Another similar relation can be derived using Eq. (12). Even though the two terms in the right-hand side are both of order ${\cal V}^{6}$, we have shown in a previous paper that the kurtosis $c_{1111}$ is significantly smaller in practice, except for central collisions Alqahtani:2025wan . The physical interpretation is that $\bar{V}_{2}$ is larger than $\sigma_{v_{n}}$ above 5% centrality Roubertie:2025qps . Neglecting $c_{0011}$, we obtain:
$\frac{MHC(v_{2}^{4},v_{3}^{2})}{v_{2}\{4\}^{2}MHC(v_{2}^{2},v_{3}^{2})}\approx-2.$ (17)
This prediction is also in fair agreement with data, as shown in Fig. 4. The experimental ratio is again systematically smaller than our prediction in absolute magnitude.
Figure 5: Same as Fig. 4 for the ratios Eqs. (18) and (20).
We now move on to the cumulants involving $v_{3}^{4}$, third and fourth lines of Eq. (13). They involve the mixed skewness $c_{0111}$, and also new, higher-order cumulants: a mixed “superskewness” CMS:2023bvg $c_{0122}$ and superkurtosis $c_{1122}$, which are cumulants of order 5 and 6 respectively. In the same way as we have neglected the kurtosis $c_{1111}$ relative to the mixed skewness $c_{0111}$ in Eq. (12), we neglect the superkurtosis $c_{1122}$ and only keep the terms involving $c_{0122}$. We then eliminate $c_{0122}$ through a linear combination, and we obtain:
$\frac{MHC(v_{2}^{4},v_{3}^{4})+2v_{2}\{4\}^{2}MHC(v_{2}^{2},v_{3}^{4})}{MHC(v_{2}^{2},v_{3}^{2})^{2}}\approx-4.$ (18)
Comparison with ALICE data is displayed in Fig. 5. The agreement with our prediction is much worse than in Fig. 4. As in Fig. 4, the ratio is smaller than our prediction in absolute magnitude. It decreases monotonically as a function of centrality, unlike the ratios in Fig. 4. Our prediction is quantitatively wrong, but its order of magnitude is correct, which is a non-trivial achievement given the simplicity of our approach.
Another similar relation can be predicted using the cumulant of order $10$ in Eq. (14), which is not yet measured:
$\frac{v_{2}\{4\}^{2}MHC(v_{2}^{6},v_{3}^{4})+6v_{2}\{4\}^{4}MHC(v_{2}^{4},v_{3}^{4})}{MHC(v_{2}^{4},v_{3}^{2})^{2}}=6.$ (19)
The advantage over Eq. (18) is one need not neglect $c_{1111}$ and $c_{1122}$, which do not enter the leading-order expressions of the cumulants involved in Eq. (19). Like Eqs. (15) and (16), Eq. (19) is a rigorous leading-order result, and we expect that it should be fairly accurate with a fine centrality binning.
Finally, we comment on the order of magnitude of $MHC(v_{2}^{2},v_{3}^{6})$, last line of Eq. (13). It is the sum of three contributions which are all of order ${\cal V}^{14}$. Only the first of these terms can be related to lower-order cumulants, whose expressions involve $c_{0111}$ and $c_{0122}$. We obtain an order of magnitude estimate by keeping only these cumulants in the right-hand side of Eqs. (12) and (13), and neglecting all other cumulants, namely, $c_{1111}$, $c_{1122}$, $c_{0133}$ and $c_{1133}$. Simple algebra then gives
$\frac{v_{2}\{4\}^{4}MHC(v_{2}^{2},v_{3}^{6})}{MHC(v_{2}^{2},v_{3}^{2})\left(MHC(v_{2}^{4},v_{3}^{4})+6v_{2}\{4\}^{2}MHC(v_{2}^{2},v_{3}^{4})\right)}\approx\frac{18}{16}.$ (20)
Agreement with data, displayed in Fig. 5, is even worse than for Eq. (18). Unlike what was observed with the previous ratios, the left-hand side is not smaller than the right-hand side for all centralities. This is not surprising, as there is no regime where the term involving $c_{0133}$ is negligible with respect to the first term.
VI Conclusions
We have derived several analytic relations between cumulants of the joint distribution between $v_{2}$ and $v_{3}$ in ultrarelativistic nucleus-nucleus collisions. For collisions at fixed impact parameter, Eqs. (15), (16) and (19) are exact in the limit of an infinitely-large system, to the extent that $v_{2}$ and $v_{3}$ are determined by linear response to anisotropies of the initial density profile. These simple and general results illustrate that cumulants are useful, not only for suppressing non-flow correlations, but also for extracting generic properties of flow fluctuations.
This simplicity arises from the general properties of local density fluctuations, which are of quantum origin. The classical fluctuations of impact parameter spoil this simplicity Roubertie:2025qps ; Alqahtani:2025wan . It is therefore essential to work with fine centrality bins. Despite the wide centrality bins used by ALICE, some of our relations are in reasonable agreement with Pb+Pb data. We expect that agreement would be significantly improved with a finer binning. If there are residual nonflow correlations, accuracy will be also improved with an increased pseudorapidity coverage.
It will be important to check whether full hydrodynamic calculations with fluctuating initial conditions, carried out at fixed impact parameter, confirm the validity of our results, which are derived under the simplifying assumption of linear hydrodynamic response. Significant deviations from linear response are observed for elliptic flow above 30% centrality Niemi:2015qia ; Noronha-Hostler:2015dbi ; Hippert:2020kde ; Giacalone:2017uqx , and their effect on mixed harmonic cumulants must be quantitatively assessed. The ALICE paper ALICE:2021adw contains full hydrodynamic calculations, done with the iEBE-VISHNU model Zhao:2017yhj . These calculations are in fair agreement with experimental results, but have large error bars. This can easily be improved. Hydrodynamic calculations of $MHC(v_{2}^{2m},v_{3}^{2q})$ with $q=1$ should require modest statistics ($q=2$ and $q=3$ are more demanding because the corresponding normalized cumulants are smaller), provided that one evaluates anisotropic flow directly on the freeze-out surface Cooper:1974mv . By contrast, the iEBE-VISHNU calculation couples the hydrodynamics to a transport calculation Bass:1998ca ; Petersen:2008dd . This is the state-of-the-art approach when it comes to comparing with experimental data, but it results in larger statistical errors, induced by the Monte Carlo sampling of hadrons. It is likely that effects of the transport phase largely cancel when taking ratios as in Sec. V. It will be also essential to carry out these hydrodynamic calculations at fixed impact parameter, which is rarely done. Impact parameter fluctuations can be studied independently Samanta:2023amp ; Alqahtani:2024ejg .
If precise hydrodynamic calculations and future experimental analyses confirm our predicions, this will be a spectacular illustration of the unreasonable effectiveness of hydrodynamics Noronha-Hostler:2015wft in describing multiparticle correlations observed in nucleus-nucleus collisions.
Acknowledgements.
M. Alqahtani acknowledges the support of the Research Mobility Program of the French Embassy in Riyadh, which helped to initiate this work.
Appendix A Moments and cumulants of $v_{2}$
The cumulant of order $2k$ of a single harmonic $v_{n}$, denoted by $c_{n}\{2k\}$, is a standard observable of heavy-ion collisions. It is obtained by setting $m=0$ or $q=0$ in Eq. (2), expanding to order $2k$ in the remaining variable, and using the definition (3). One obtains the well-known relations
$c_{n}\{2\}$ $=v_{n}\{2\}^{2}=\langle v_{n}^{2}\rangle$
$c_{n}\{4\}$ $=-v_{n}\{4\}^{4}=\langle v_{n}^{4}\rangle-2\langle v_{n}^{2}\rangle^{2}$
$c_{n}\{6\}$ $=4v_{n}\{6\}^{6}=\langle v_{n}^{6}\rangle-9\langle v_{n}^{4}\rangle\langle v_{n}^{2}\rangle+12\langle v_{n}^{2}\rangle^{3}.$ (21)
Using Eqs. (2), (3), (9) and (11), one shows that for $n=2$, cumulants of order 4 and higher are solely determined by the mean elliptic flow in the reaction plane, $v_{2}\{4\}=v_{2}\{6\}=\bar{V}_{2}$, up to relative corrections of order ${\cal V}^{2}$ (at the sub-percent level). These corrections originate mostly from the skewness of elliptic flow fluctuations Giacalone:2016eyu ; CMS:2017glf ; ALICE:2018rtz ; CMS:2023bvg , which is $c_{1200}$ in our notation Roubertie:2025qps .
Inverting Eqs. (21), one expresses the moments in terms of cumulants:
$\langle v_{n}^{2}\rangle$ $=v_{n}\{2\}^{2}$
$\langle v_{n}^{4}\rangle$ $=2v_{n}\{2\}^{4}-v_{n}\{4\}^{4}$
$\langle v_{n}^{6}\rangle$ $=6v_{n}\{2\}^{6}-9v_{n}\{2\}^{2}v_{n}\{4\}^{4}+4v_{n}\{6\}^{6}.$ (22)
We use these equations to evaluate the denominators in Eq. (5). The cumulants in the right-hand side of Eq. (22) are measured in the same experiment with the same centrality bins ALICE:2016ccg . We neglect $v_{3}\{4\}$ and $v_{3}\{6\}$, which are not provided in this reference, and whose contribution is known to be small ATLAS:2019peb . For the 0-5% centrality bin, $v_{2}\{4\}$ and $v_{2}\{6\}$ are undefined (the corresponding cumulants change sign STAR:2015mki ; ATLAS:2019peb ; Alqahtani:2024ejg ) and we set them to zero.

### v2

@@
 Abstract
-The ALICE Collaboration has analyzed a number of cumulants mixing elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), involving up to $8$ particles, in Pb+Pb collisions at the LHC.
+The ALICE and CMS Collaborations have analyzed a number of cumulants mixing elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), involving up to $8$ particles, in Pb+Pb collisions at the LHC.
 We unravel an unexpected simplicity in these complex mathematical quantities for collisions at fixed impact parameter.
@@
 We derive simple analytic relations between cumulants of different orders on this basis.
-Some of these relations are in reasonable agreement with existing data.
-We postulate that agreement will be much improved if the analysis is repeated with a finer centrality binning and a larger pseudorapidity acceptance.
+These relations are in good agreement with recent data from the CMS Collaboration.
+We argue that agreement will be further improved if the analysis is repeated with a finer centrality binning.
+We make quantitative predictions for cumulants of order 10 which have not yet been analyzed.
 I Introduction
@@
 This turns out to be the most robust aspect of global theory-to-data comparison JETSCAPE:2020mzn .
-In this paper, we unravel the information contained in cumulants of azimuthal correlations involving both elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), which are called “mixed harmonic cumulants” (MHC) ALICE:2021adw .
+In this paper, we unravel the information contained in cumulants of azimuthal correlations involving both elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), which are called “mixed harmonic cumulants” (MHC) ALICE:2021adw ; CMS:2025opi .
 $v_{2}$ and $v_{3}$ are the largest Fourier harmonics of anisotropic flow ALICE:2011ab .
@@
 The lowest-order MHC is a 4-particle cumulant introduced by Bilandzic et al.
-Bilandzic:2013kga , which has been measured in Pb+Pb collisions at the LHC by ALICE ALICE:2016kpq and ATLAS ATLAS:2019peb .
+Bilandzic:2013kga , which has been measured in Pb+Pb collisions at the LHC ALICE:2016kpq ; CMS:2025opi ; ATLAS:2019peb .
 It represents the linear correlation between $v_{2}^{2}$ and $v_{3}^{2}$, which is negative except in ultra-central collisions.
 We have recently shown Alqahtani:2025wan that it is driven by the correlation between the elliptic flow in the reaction plane, $v_{2,x}$, and $v_{3}^{2}$.
-Our goal is to extend this study to higher-order cumulants, of 6 and 8 particles, which have subsequently been measured ALICE:2021adw .
+Our goal is to extend this study to higher-order cumulants, of 6 and 8 particles, which have subsequently been measured ALICE:2021adw ; CMS:2025opi .
 Throughout this paper, we assume that nonflow correlations are negligible, so that particles in each event are emitted independently according to an underlying probability distribution Luzum:2011mm ; Ollitrault:2023wjk .
@@
 In Sec.
-II, we define the cumulants of $v_{2}$ and $v_{3}$ measured by ALICE using the formalism of generating functions and we briefly discuss the data.
+II, we define the cumulants of $v_{2}$ and $v_{3}$ using the formalism of generating functions and we compare data from several analyses in Pb+Pb collisions at the LHC.
 In Sec.
@@
 IV, we express the experimental cumulants as a function of the cumulants in the intrinsic frame.
-We explain the orders of magnitude of cumulants measured by ALICE.
 In Sec.
@@
 II Cumulants of flow fluctuations in the laboratory frame
+II.1 Definitions
 The quantities that can be measured experimentally are moments of the joint distribution of $V_{2}$ and $V_{3}$ Bhalerao:2011yg ; Bhalerao:2014xra .
-The cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$ analyzed by ALICE ALICE:2021adw , where $m$ and $q$ are positive integers, are combinations of these moments Bilandzic:2010jr ; Bilandzic:2013kga .
+The cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$ analyzed by ALICE ALICE:2021adw and CMS CMS:2025opi , where $m$ and $q$ are positive integers, are combinations of these moments Bilandzic:2010jr ; Bilandzic:2013kga .
 [^1] The ALICE collaboration has subsequently analyzed a different set of “cumulants” which are different combination of moments ALICE:2023lwx .
@@
 We do not repeat them because they are lengthy and will not be needed in this work.
-[^3] Note that the six MHC cumulants analyzed by ALICE only involve the magnitudes, not the phases of $V_{2}$ and $V_{3}$.
+[^3] Note that the six MHC cumulants analyzed by ALICE and CMS only involve the magnitudes, not the phases of $V_{2}$ and $V_{3}$.
 The phases only enter at the next order $(m,q)=(3,2)$, a cumulant of order 10 whose expression involves moments of the type $\langle V_{2}^{3}(V_{3}^{*})^{2}\rangle$, and which has not yet been analyzed.
-Figure 1: Comparison between ALICE ALICE:2021adw and ATLAS ATLAS:2019peb data on $nMHC(v_{2}^{2},v_{3}^{2})$ in Pb+Pb collisions at 5.02 TeV per nucleon pair, as a function of the collision centrality.
-ATLAS denotes this quantity by $nsc_{2,3}\{4\}$.
-ALICE normalizes the cumulants as follows:
+The cumulants are usually normalized as follows:
 $nMHC(v_{2}^{2m},v_{3}^{2q})\equiv\frac{MHC(v_{2}^{2m},v_{3}^{2q})}{\langle v_{2}^{2m}\rangle\langle v_{3}^{2q}\rangle}.$ (5)
-This normalization suppresses the sensitivity to kinematic cuts, and also provides an intuitive, dimensionless measure of the strength of the correlation.
-A value of order unity would imply a strong correlation between $v_{2}$ and $v_{3}$.
-But the measured values of $nMHC(v_{2}^{2m},v_{3}^{2q})$ are, in absolute magnitude, smaller than unity for all centralities.
+This normalization suppresses the sensitivity to kinematic cuts: If $v_{n}$ is proportional to the initial anisotropy $\varepsilon_{n}$, the dependence on kinematic cuts can only be in the proportionality factor (the linear response coefficient) which cancels in the ratio (5).
+The normalized coefficient provides an intuitive dimensionless measure of the strength of the correlation.
+A value of order unity implies a strong correlation between $v_{2}$ and $v_{3}$.
+Figure 1: Centrality dependence of $nMHC(v_{2}^{2},v_{3}^{2})$ in Pb-Pb collisions at 5.02 TeV: Comparison between 2021 ALICE:2021adw and 2024 ALICE:2024nqd ALICE results, ATLAS ATLAS:2019peb and CMS CMS:2025opi .
+Note that this quantity is denoted by $NSC(3,2)$ in the 2024 ALICE paper, and by $nsc_{2,3}\{4\}$ in the ATLAS paper.
+II.2 Experimental data
+There have been several analyses of the lowest-order cumulant $nMHC(v_{2}^{2},v_{3}^{2})$ in Pb+Pb collisions at the LHC, which we first compare.
+We then compare ALICE and CMS results for the higher-order cumulants.
+The normalization (5) facilitates the comparison between experiments, because of the reduced sensitivity to kinematic cuts.
+Fig.
+1 displays four sets of experimental results for $nMHC(v_{2}^{2},v_{3}^{2})$ obtained by the ALICE Collaboration in 2021 ALICE:2021adw and 2024 ALICE:2024nqd , by ATLAS ATLAS:2019peb , and recently by CMS CMS:2025opi .
+One notices that the 2024 ALICE results differ significantly from the 2021 results between 10% and 30% centrality.
+The 2024 analysis is done in narrower bins (5% as opposed to 10%).
+The sensitivity of this specific observable to the width of the centrality bins has already been pointed out Gardim:2016nrr , and narrower bins are preferred.
+Results from ATLAS, CMS, and ALICE 2024 are in fair agreement.
+The origin of the residual discrepancies is unknown.
+It could be due to the kinematic cuts, which are different for the four analyses.
+Note that the dependence on $p_{T}$ cuts cannot be explained within the simple picture of linear hydrodynamic response.
+The negative sign of $nMHC(v_{2}^{2},v_{3}^{2})$ and the centrality dependence can be ascribed to non-Gaussian fluctuations in the initial state Alqahtani:2025wan , as will be recalled in Sec.
+IV.
+Figure 2: Comparison between ALICE ALICE:2021adw and CMS CMS:2025opi results for higher-order normalized mixed harmonic cumulants.
+In this paper, we study higher-order cumulants which were measured in by ALICE in 2021 ALICE:2021adw and more recently by CMS CMS:2025opi .
+The two sets of results are compared in Fig.
+2 as a function of the collision centrality.
+The normalized cumulants are, in absolute magnitude, smaller than unity, which means that correlations between $v_{2}$ and $v_{3}$ are not strong.
 [^4] By contrast, $nMHC(v_{2}^{2},v_{4}^{2})$ reaches $1$ in peripheral collisions ATLAS:2019peb ; ALICE:2021adw , which is a natural consequence of the non-linear coupling between $v_{2}$ and $v_{4}$ Teaney:2012ke .
-They increase somewhat with $m$, and decrease strongly with $q$, which will be explained in Sec.
+The sign of $nMHC(v_{2}^{2m},v_{3}^{2q})$ alternates as $m$ or $q$ increases.
+In absolute magnitude, it increases somewhat with $m$, and decreases strongly with $q$.
+These features will be explained in Sec.
 IV.
-The normalization (5) also facilitates the comparison with other experiments and theory calculations.
-A comparison between ALICE and ATLAS ATLAS:2019peb data for the lowest-order cumulant $nMHC(v_{2}^{2},v_{3}^{2})$ is displayed in Fig.
-1.
-There are sizable differences between the two experiments, whose origin is unknown.
-One possible explanation is the wider centrality bins used by ALICE.
-However, one would typically expect wider bins to increase the value of $nMHC(v_{2}^{2},v_{3}^{2})$ Gardim:2016nrr , and ALICE is below ATLAS for most centralities.
-Interestingly, ATLAS observes a variation of $nMHC(v_{2}^{2},v_{3}^{2})$ going in the same direction as ALICE (down for centralities $40\%$, up for centralities $40\%$) when only particles with $p_{T}>2$ GeV/c/c are included.
-Since these high-$p_{T}$ particles are likely to have sizable nonflow correlations from jet production, it is tempting to postulate that the difference between ATLAS and ALICE may be due to larger nonflow effects in ALICE, due to the smaller acceptance in pseudorapidity.
-We consider it likely that ATLAS results are more precise than ALICE results, and we will take this as an excuse for not understanding precisely the ALICE results on higher-order cumulants in Sec.
-V.
-Figure 2: Mixed harmonic cumulants as a function of centrality for Pb+Pb collisions at 5.02 TeV per nucleon pair.
-The normalized cumulants (5) are taken from Ref.
-ALICE:2021adw , and the moments in the denominator are evaluated using earlier ALICE data ALICE:2016ccg and formulas in Appendix A.
-For the negative cumulants, we change the sign before taking the logarithm.
-$MHC(v_{2}^{4},v_{3}^{4})$ and $MHC(v_{2}^{2},v_{3}^{6})$ are positive for the most central bins, and the three corresponding data points are circled.
+As already noticed for the lowest-order cumulant in Fig.
+1, there are sizable differences between ALICE and CMS results.
+CMS uses 5% centrality bins between 10% and 40% centrality, as opposed to 10% bins for ALICE, which may explain some of the discrepancies in this range.
+The CMS cumulants increase monotonically with the centrality fraction in absolute magnitude in agreement with hydrodynamics predictions shown in ALICE:2021adw ; CMS:2025opi , while the ALICE cumulants reach a maximum and then decrease, and we do not know the origin of this difference.
+Figure 3: Mixed harmonic cumulants as a function of centrality for Pb+Pb collisions at 5.02 TeV per nucleon pair, using CMS data CMS:2025opi for the normalized cumulants and evaluating the denominator of (5) with formulas in Appendix A.
 For our analysis, we will need the un-normalized cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$.
-In order to compute them, we evaluate the moments appearing in the denominator of Eq.
+In order to obtain them from the normalized cumulants, we evaluate the moments appearing in the denominator of Eq.
 (5) using standard formulas which are recalled in Appendix A.
-Results are displayed in Fig.
+For the sake of illustration, CMS results are displayed in Fig.
 3 as a function of the collision centrality.
-The lowest-order cumulant defined by Eq.
-(4) is negative, as shown in Fig.
-1, and the sign alternates as $m$ or $q$ increases, as illustrated by the legend of the figure.
+One sees that they span ten orders of magnitude, a point to which we come below in Sec.
+IV.
 III Cumulants of flow fluctuations in the intrinsic frame
@@
 We call this the “intrinsic frame” Alqahtani:2024ejg ; Roubertie:2025qps , as illustrated in Fig.
-3.
+4.
 More detailed information is available in the intrinsic frame than in the laboratory frame, in which the direction of impact parameter is unknown.
@@
 Two weaker symmetries remain: Symmetry with respect to the reaction plane, $\varphi\to-\varphi$ (Fig.
-3), implies that $c_{mpqr}$ is real and $c_{pmrq}=c_{mpqr}$.
+4), implies that $c_{mpqr}$ is real and $c_{pmrq}=c_{mpqr}$.
 Symmetry between target and projectile, $\varphi\to\varphi+\pi$, further implies that $c_{mpqr}$ vanishes for odd $q+r$.
@@
 $\sigma_{v_{3}}^{2}$ $\equiv c_{0011}.$ (8)
-The central limit theorem implies that only cumulants of order $1$ and $2$ are kept, which means that the probability distribution of $V_{2}$ and $V_{3}$ is Gaussian in the intrinsic frame Voloshin:2007pc .
-Since none of these cumulants mix $V_{2}$ and $V_{3}$, they are uncorrelated in the Gaussian limit Alqahtani:2025wan .
-Therefore, the correlations studied in this paper are all driven by non-Gaussian properties of the distribution, so that we shall not neglect cumulants of order $3$ and higher.
+If fluctuations in the intrinsic frame are Gaussian Voloshin:2007pc , all cumulants vanish beyond order $2$.
+Now, the cumulants of order $1$ and $2$ listed above do not mix $V_{2}$ and $V_{3}$.
+Therefore, $V_{2}$ and $V_{3}$ are uncorrelated in the Gaussian limit Alqahtani:2025wan , and the correlations studied in this paper are all driven by non-Gaussian properties of the distribution.
+In order to model these non-Gaussianities, we need to take into account cumulants of order $3$ and higher.
 [^7] Non-Gaussian fluctuations are also responsible for the non-zero $v_{3}\{4\}$ seen in Pb+Pb ALICE:2011ab and Xe+Xe CMS:2019cyz ; CMS:2025xtt collisions , the hierarchy of higher-order cumulants of $v_{2}$ ($v_{2}\{4\}$, $v_{2}\{6\}$, $v_{2}\{8\}$) in p+Pb collisions CMS:2015yux ; CMS:2019wiy , as well as the splitting between $v_{2}\{4\}$ and $v_{2}\{6\}$ in mid-central Pb+Pb collisions Giacalone:2016eyu ; CMS:2017glf ; ALICE:2018rtz ; CMS:2023bvg .
@@
 For simplicity, we neglect centrality fluctuations, and we assume that all events in a centrality class have the same impact parameter.
-The ALICE experiment uses wide centrality bins where this approximation may not be good enough, but this can easily be improved in future analyses.
-Centrality fluctuations always matter in the 0–5% centrality window Alqahtani:2024ejg , but we have found that their effect on correlations between $v_{2}$ and $v_{3}$ is modest beyond 5% centrality Alqahtani:2025wan , provided that the analysis uses narrow centrality bins ATLAS:2019peb , and we hope that the same conclusion applies to the higher-order cumulants studied in this paper.
+Impact parameter fluctuations have large effects for the most central collisions Alqahtani:2024ejg , but we have found that their effect on correlations between $v_{2}$ and $v_{3}$ is modest beyond 5% centrality Alqahtani:2025wan , provided that the analysis uses narrow centrality bins ATLAS:2019peb .
+The ALICE experiment uses wide centrality bins and this has a significant effect on the results, as discussed in Sec.
+II.2, but this can easily be improved in future analyses.
 If the impact parameter is constant, different events in the intrinsic frame differ only by local fluctuations in the initial density profile, which are of quantum origin, as illustrated in Fig.
-3.
+4.
 The resulting fluctuations of the complex anisotropies $\varepsilon_{2}$ and $\varepsilon_{3}$ Qiu:2011iv , which generate $V_{2}$ and $V_{3}$ through linear hydrodynamic response, originate from a large number $N$ of independent local fluctuations.
@@
 $MHC(v_{2}^{2},v_{3}^{2})$ $=c_{1000}c_{0111}+c_{0100}c_{1011}+c_{1111}$
-$=2\bar{V}_{2}c_{0111}+c_{1111}$ (12)
+$=2\bar{V}_{2}c_{0111}+c_{1111}.$ (12)
 where, in the last equality, we have introduced $\bar{V}_{2}$ defined by Eq.
@@
 (10).
-We now list the expressions of the higher-order cumulants measured by ALICE ALICE:2021adw , which we truncate by keeping only the leading terms in the power-series expansion:
+We now list the expressions of the higher-order cumulants measured by ALICE and CMS ALICE:2021adw ; CMS:2025opi , which we truncate by keeping only the leading terms in the power-series expansion:
 $MHC(v_{2}^{4},v_{3}^{2})$ $=-4\,\bar{V}_{2}^{3}\,c_{0111}$
@@
 These orders of magnitude are reflected in the hierarchy observed in Fig.
-2.
+3.
 In particular, they explain why $MHC(v_{2}^{6},v_{3}^{2})$ and $MHC(v_{2}^{2},v_{3}^{4})$ are of comparable magnitude, despite being cumulants of different orders (8 and 6 respectively).
 Since $v_{2}$ and $v_{3}$ are both of order $\cal V$}, the normalized symmetric cumulant (5) is of order ${\cal V}^{2q}$.
-This explains why the magnitude of $nMHC(v_{2}^{2m},v_{3}^{2q})$ decreases strongly as $q$ increases, as pointed out in Sec.
-II.
+This explains why the magnitude of $nMHC(v_{2}^{2m},v_{3}^{2q})$ decreases strongly as $q$ increases, as seen in Fig.
+2.
 We finally provide leading-order expressions for the cumulants of order 10, which have not yet been analyzed, but could easily be measured:
@@
 A first comment on the expressions (12), (13) and (14) is that they only involve the mean elliptic flow in the reaction plane, $\bar{V}_{2}$, and mixed cumulants of order $\geq 3$, which quantify non-Gaussian fluctuations.
+Thus the correlations between $v_{2}$ and $v_{3}$ originate from non-Gaussian fluctuations in the initial state Alqahtani:2025wan .
 A second remark is that for a given value of $q$, $MHC(v_{2}^{2},v_{3}^{2q})$ contains more terms than higher-order cumulants $MHC(v_{2}^{2m},v_{3}^{2q})$ with $m\geq 2$, whose expressions are somewhat simpler.
@@
 This can be related to the well-known fact that cumulants of $v_{2}$ of order 4 and higher only depend on $\bar{V}_{2}$ to leading order in $\cal V$}, as recalled in Appendix A.
-In addition, there are overall numerical factors which increase rapidly as a function of $m$ (22, 44, $24$ for $m=1,2,3$).
-These numerical factors explain why the normalized cumulants $nMHC(v_{2}^{2m},v_{3}^{2q})$ increase as a function of $m$ for fixed $q$, in absolute magnitude ALICE:2021adw .
+In addition, there are overall numerical factors which increase rapidly as a function of $m$ (22, 4-4, $24$, 264-264 for $m=1,2,3,4$).
+These numerical factors explain why the normalized cumulant $nMHC(v_{2}^{2m},v_{3}^{2q})$ changes sign and increases in absolute magnitude as $m$ increases, as seen in Fig.
+2.
 V Relations between cumulants of different orders
 Figure 5: Ratios in Eqs.
+(15) and (17) as a function of the collision centrality in Pb+Pb collisions at 5.02 TeV per nucleon pair.
+Open symbols are ALICE data, where the mixed cumulants are taken from Ref.
+ALICE:2021adw and $v_{2}\{4\}$ from Ref.
+ALICE:2016ccg .
+Closed symbols are CMS data CMS:2025opi .
+Horizontal lines are our theory predictions, Eqs.
 (15) and (17).
-Symbols are ALICE data, where the mixed cumulants are taken from Ref.
-ALICE:2021adw and $v_{2}\{4\}$ from Ref.
-ALICE:2016ccg , as a function of the collision centrality in Pb+Pb collisions at 5.02 TeV per nucleon pair.
-Horizontal lines are our theory predictions.
 Eqs.
@@
 A).
-Since $v_{2}\{4\}$ is only measured above 5% centrality ALICE:2016ccg , we exclude the most central bin 0-5% from now on.
+ALICE only measures $v_{2}\{4\}$ above 5% centrality ALICE:2016ccg so that we exclude the most central bin 0-5%.
 The remaining intrinsic cumulants can be eliminated by combining the information from several measured cumulants. Taking the ratio between the first two lines of Eq. (13), for instance, one obtains:
 $\frac{MHC(v_{2}^{6},v_{3}^{2})}{v_{2}\{4\}^{2}MHC(v_{2}^{4},v_{3}^{2})}=-6.$ (15)
-This prediction is tested against ALICE data in Fig.
-4.
-The experimental ratio in the left-hand side is in rough agreement with the predicted value in the right-hand side, but somewhat smaller in absolute magnitude.
+Fig. 5 shows that CMS data agree very well with this equation, except for the most central bin, where impact parameter fluctuations may have a sizable effect. Agreement with ALICE data is poorer.
 Similarly, using Eqs. (13) and (14), we predict:
@@
 This could easily be checked experimentally, as increasing the order in $v_{2}$ does not significantly increase errors.
-Eqs.
-(15) and (16) are rigorous mathematical results to leading order in $\cal V$}.
+Equations (15) and (16) are rigorous mathematical results to leading order in $\cal V$}.
 They generalize the well-known identities $v_{2}\{4\}=v_{2}\{6\}=v_{2}\{8\}$ (Appendix A) to mixed cumulants.
-We therefore expect that their accuracy is comparable, at the percent level in Pb+Pb collisions at LHC energies CMS:2017glf ; ALICE:2018rtz ; CMS:2023bvg .
-The discrepancy between our prediction and data is an order of magnitude larger than expected in Fig.
-4.
-We predict that agreement will be much improved if the analysis is repeated with finer centrality bins.
+We therefore expect that their accuracy is comparable, at the percent level in Pb+Pb collisions at LHC energies CMS:2017glf ; ALICE:2018rtz ; CMS:2023bvg , provided that the analysis is carried out in fine centrality bins.
+[^11] The intrinsic centrality resolution at the LHC is of order 2% Das:2017ned .
+This is the typical order of magnitude of the optimal width of the centrality binning.
 Another similar relation can be derived using Eq.
@@
 $\frac{MHC(v_{2}^{4},v_{3}^{2})}{v_{2}\{4\}^{2}MHC(v_{2}^{2},v_{3}^{2})}\approx-2.$ (17)
-This prediction is also in fair agreement with data, as shown in Fig. 4. The experimental ratio is again systematically smaller than our prediction in absolute magnitude.
+This prediction is also in good agreement with data, as shown in Fig.
+5.
+Eqs.
+(15), (16), and (17) show that as one increases $m$, the change in $MHC(v_{2}^{2m},v_{3}^{2})$ is solely determined by the mean elliptic flow in the reaction plane $\bar{V}_{2}\simeq v_{2}\{4\}$.
+This property also holds for higher orders in $v_{3}$, which we now study.
 Figure 6: Same as Fig. 5 for the ratios Eqs. (18) and (20).
@@
 $\frac{MHC(v_{2}^{4},v_{3}^{4})+2v_{2}\{4\}^{2}MHC(v_{2}^{2},v_{3}^{4})}{MHC(v_{2}^{2},v_{3}^{2})^{2}}\approx-4.$ (18)
-Comparison with ALICE data is displayed in Fig.
-5.
-The agreement with our prediction is much worse than in Fig.
-4.
-As in Fig.
-4, the ratio is smaller than our prediction in absolute magnitude.
-It decreases monotonically as a function of centrality, unlike the ratios in Fig.
-4.
-Our prediction is quantitatively wrong, but its order of magnitude is correct, which is a non-trivial achievement given the simplicity of our approach.
+Comparison with experimental data is displayed in Fig. 6. Our prediction agrees well with CMS data, not with ALICE data.
 Another similar relation can be predicted using the cumulant of order $10$ in Eq. (14), which is not yet measured:
@@
 $\frac{v_{2}\{4\}^{4}MHC(v_{2}^{2},v_{3}^{6})}{MHC(v_{2}^{2},v_{3}^{2})\left(MHC(v_{2}^{4},v_{3}^{4})+6v_{2}\{4\}^{2}MHC(v_{2}^{2},v_{3}^{4})\right)}\approx\frac{18}{16}.$ (20)
-Agreement with data, displayed in Fig.
-5, is even worse than for Eq.
-(18).
-Unlike what was observed with the previous ratios, the left-hand side is not smaller than the right-hand side for all centralities.
-This is not surprising, as there is no regime where the term involving $c_{0133}$ is negligible with respect to the first term.
+This prediction roughly matches CMS data in terms of order of magnitude (Fig.
+6).
+Here, we do not expect any significant improvement with a finer centrality binning, as we have neglected the term involving $c_{0133}$ which is generically of the same order of magnitude as the other terms.
 VI Conclusions
@@
 It is therefore essential to work with fine centrality bins.
-Despite the wide centrality bins used by ALICE, some of our relations are in reasonable agreement with Pb+Pb data.
-We expect that agreement would be significantly improved with a finer binning.
-If there are residual nonflow correlations, accuracy will be also improved with an increased pseudorapidity coverage.
+The wide centrality bins implemented by ALICE partly explain the differences between data and our predictions.
+The recent CMS data are in much better agreement with our calculation.
+This illustrates that the effect of impact parameter fluctuations is modest.
+We expect that agreement would be further improved with a finer centrality binning.
+In addition, we have made a number of quantitative predictions for cumulants of order 10 (Eqs.(16) and (19)), which could be analyzed with existing data.
 It will be important to check whether full hydrodynamic calculations with fluctuating initial conditions, carried out at fixed impact parameter, confirm the validity of our results, which are derived under the simplifying assumption of linear hydrodynamic response.
 Significant deviations from linear response are observed for elliptic flow above 30% centrality Niemi:2015qia ; Noronha-Hostler:2015dbi ; Hippert:2020kde ; Giacalone:2017uqx , and their effect on mixed harmonic cumulants must be quantitatively assessed.
-The ALICE paper ALICE:2021adw contains full hydrodynamic calculations, done with the iEBE-VISHNU model Zhao:2017yhj .
+The ALICE ALICE:2021adw and CMS CMS:2025opi include comparisons of data with full hydrodynamic calculations, done with the iEBE-VISHNU model Zhao:2017yhj and with the IP-GLASMA+MUSIC+URQMD framework McDonald:2017ayb ; Gale:2013da ; Bass:1998ca .
 These calculations are in fair agreement with experimental results, but have large error bars.
@@
 Hydrodynamic calculations of $MHC(v_{2}^{2m},v_{3}^{2q})$ with $q=1$ should require modest statistics ($q=2$ and $q=3$ are more demanding because the corresponding normalized cumulants are smaller), provided that one evaluates anisotropic flow directly on the freeze-out surface Cooper:1974mv .
-By contrast, the iEBE-VISHNU calculation couples the hydrodynamics to a transport calculation Bass:1998ca ; Petersen:2008dd .
+By contrast, both iEBE-VISHNU and IP-GLASMA+MUSIC +URQMD couple the hydrodynamics to a transport calculation Bass:1998ca ; Petersen:2008dd .
 This is the state-of-the-art approach when it comes to comparing with experimental data, but it results in larger statistical errors, induced by the Monte Carlo sampling of hadrons.
@@
 It will be also essential to carry out these hydrodynamic calculations at fixed impact parameter, which is rarely done.
-Impact parameter fluctuations can be studied independently Samanta:2023amp ; Alqahtani:2024ejg .
-If precise hydrodynamic calculations and future experimental analyses confirm our predicions, this will be a spectacular illustration of the unreasonable effectiveness of hydrodynamics Noronha-Hostler:2015wft in describing multiparticle correlations observed in nucleus-nucleus collisions.
+Impact parameter fluctuations can be studied independently Samanta:2023amp ; Alqahtani:2024ejg , and we have argued that their effects should be small above 5-10% centrality.
 Acknowledgements.
-M. Alqahtani acknowledges the support of the Research Mobility Program of the French Embassy in Riyadh, which helped to initiate this work.
+We thank Ante Bilandzic and You Zhou for discussions about the ALICE results.
+M.
+Alqahtani acknowledges the support of the Research Mobility Program of the French Embassy in Riyadh, which helped to initiate this work.
 Appendix A Moments and cumulants of $v_{2}$
@@
 $\langle v_{n}^{6}\rangle$ $=6v_{n}\{2\}^{6}-9v_{n}\{2\}^{2}v_{n}\{4\}^{4}+4v_{n}\{6\}^{6}.$ (22)
-We use these equations to evaluate the denominators in Eq.
+The cumulants in the right-hand side of Eq.
+(22) are taken from each experiment and used to evaluate the denominators in Eq.
 (5).
-The cumulants in the right-hand side of Eq.
-(22) are measured in the same experiment with the same centrality bins ALICE:2016ccg .
+For ALICE results, they are measured with the same centrality bins ALICE:2016ccg .
 We neglect $v_{3}\{4\}$ and $v_{3}\{6\}$, which are not provided in this reference, and whose contribution is known to be small ATLAS:2019peb .
 For the 0-5% centrality bin, $v_{2}\{4\}$ and $v_{2}\{6\}$ are undefined (the corresponding cumulants change sign STAR:2015mki ; ATLAS:2019peb ; Alqahtani:2024ejg ) and we set them to zero.
+On the other hand, for CMS results they are taken from CMS:2025opi with finer centrality bins.
+We note here that $v_{3}\{4\}$ is provided in this CMS analysis.